# Unifying Bayesian Flow Networks and Diffusion Models through Stochastic Differential Equations

### v1

@@ 4.1 Formulating BFN on Continuous Data as SDEs
The time $t$ is truncated by $1-\eta$ in Theorem 4.1 for two reasons. On one hand, the reverse-time SDE derived later (see Eq. (20)) is ill-defined at $t=1$ since the distribution of ${\bm{\mu}}$ collapses to a Dirac distribution whose score tends to infinity. On the other hand, it is convenient to satisfy certain regularity conditions for the uniqueness of the solution in Theorem 4.1, as detailed in the proof. As $\eta$ is small (e.g., $10^{-3}\sim 10^{-5}$) in our implementation, the effect of truncation is negligible. The exact distribution of ${\bm{\mu}}(1-\eta)$ is unknown, we approximate it by an isotropic Gaussian with a zero mean and small variance (see details in Sec. 6).
@@ 4.4 Probability Flow ODE and Faster Sampling
where $h_{i}=\lambda(t_{i})-\lambda(t_{i-1})$. We refer the readers to Appendix A.3 for higher-order solvers of both ODE and SDE. [^4] A more straightforward way to get BFN-Solver on continuous data is to treat BFN as a DM with a special noise schedule $\alpha_{t}=\gamma_{t}$ and $\sigma^{2}_{t}=\gamma_{t}(1-\gamma_{t})$. However, it is infeasible on discrete data. Therefore, we use a slightly complex yet coherent way to derive BFN-Solver throughout the paper.
@@ 5.3 Sampling as Discretizing the Reverse-time SDE
The role of the categorical sampling step is still unclear in theory. However, experiments in Fig. 6 (Sec. 6.3) reveal that removing the categorical sampling step leads to consistently better performance in fewer than $50$ NFEs, and almost the same results otherwise.
@@ 6.1 Experimental Settings
Datasets. For continuous data, the model is trained on the CIFAR-10 (Krizhevsky et al., 2009) dataset which contain 50K training images. For discrete data, the model is trained on the text8 (Mahoney, 2011) dataset which contains 90M consecutive characters, each character is a lower Latin letter ‘a’-‘z’ or the ‘__\__’ whitespace token, giving a class number of 27. Each sample is a sequence of 256 characters.
@@ 6.3 Fast Sampling on Discrete Data
We compare our proposed fast sampling methods with the origin BFN discrete sampler in this section.

### v2

@@ 4.1 Formulating BFN on Continuous Data as SDEs
The time $t$ is truncated by $1-\eta$ in Theorem 4.1 for two reasons. On one hand, the reverse-time SDE derived later (see Eq. (20)) is ill-defined at $t=1$ since the distribution of ${\bm{\mu}}$ collapses to a Dirac distribution whose score tends to infinity. On the other hand, it is convenient to satisfy certain regularity conditions for the uniqueness of the solution in Theorem 4.1, as detailed in the proof. The exact distribution of ${\bm{\mu}}(1-\eta)$ is unknown. We approximate it by an isotropic Gaussian with zero mean and small variance, as $p({\bm{\mu}}(t))$ tends towards a Dirac delta function when $t\to 1$. (see details in Sec. 6). As $\eta$ is small (e.g., $10^{-3}\sim 10^{-5}$) in our implementation, the effect of truncation is negligible.
@@ 4.4 Probability Flow ODE and Faster Sampling
where $h_{i}=\lambda(t_{i})-\lambda(t_{i-1})$. BFN-Solver++ shares the same spirit with BFN-Solver and the difference is that BFN-Solver++ considers data prediction instead of noise prediction. We refer the readers to Appendix A.3 for higher-order solvers of both ODE and SDE. [^4] A more straightforward way to get BFN-Solver on continuous data is to treat BFN as a DM with a special noise schedule $\alpha_{t}=\gamma_{t}$ and $\sigma^{2}_{t}=\gamma_{t}(1-\gamma_{t})$. However, it is infeasible on discrete data. Therefore, we use a slightly complex yet coherent way to derive BFN-Solver throughout the paper.
@@ 5.3 Sampling as Discretizing the Reverse-time SDE
The role of the categorical sampling step is still unclear in theory. However, experiments in Fig. 6 (Sec. 6.3) reveal that removing the categorical sampling step leads to consistently better performance in fewer than $50$ NFEs, and almost the same results otherwise. We provide preliminary analyses of it in Appendix C.4.
@@ 6.1 Experimental Settings
Datasets. For continuous data, the model is trained on the CIFAR-10 (Krizhevsky & Hinton, 2009) dataset which contain 50K training images. For discrete data, the model is trained on the text8 (Mahoney, 2011) dataset which contains 90M consecutive characters, each character is a lower Latin letter ‘a’-‘z’ or the ‘__\__’ whitespace token, giving a class number of 27. Each sample is a sequence of 256 characters.
@@ 6.3 Fast Sampling on Discrete Data
We compare our proposed fast sampling methods with the origin BFN discrete sampler and other representative baselines in this section.
@@
We compare the efficiency of BFN-Solver2 with representative baselines including D3PM (Austin et al., 2023) and OAARDM (Hoogeboom et al., 2022). As summarized in Table 2, BFN-Solver2 outperforms all baselines with both $20$ and $50$ steps indirectly. Notably, evaluating the exact likelihood by the proposed samplers requires a nontrivial extension (Song et al., 2021). We use the original BFN sampler, evaluated under both metrics, as a bridge for an indirect comparison. For instance, according to Table 2, BFN-Solver2 ($20$ steps, SA $0.84$) is better than the original BFN sampler ($25$ steps, SA $0.80$, BPC $1.52$) and better than OA-ARDM ($20$ steps, BPC $1.52$).
[…]
Methods NFE BPC $\downarrow$ SA $\uparrow$
BFN-Solver2 (Ours) 50 - 0.85
BFN 50 1.47 0.84
D3PM (Best) 50 1.52 -
BFN 25 1.52 0.80
BFN-Solver2 (Ours) 20 - 0.84
BFN 20 1.56 0.79
BFN-Solver2 (Ours) 20 1.56 -
OA-ARDM 20 1.52 -
@@ 7 Conclusion
Acknowledgements
This work was supported by NSF of China (No. 62076145), Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Renmin University of China, the Fundamental Research Funds for the Central Universities, the Research Funds of Renmin University of China (No. 22XNKJ13), and the Ant Group Research Fund. C. Li was also sponsored by Beijing Nova Program (No. 20220484044). The work was partially done at the Engineering Research Center of Next-Generation IntelligentSearch and Recommendation, Ministry of Education.
@@ C.3 User Study for Text Generation
C.4 Ablation of Categorical Sampling Step in BFN Sampler on Discrete Data
In this section, we present the detailed results of Sec. 6.2 in Table 3. We explicitly analyze the sampling dynamics of the original BFN sampler with and without CS. We generate two sets of trajectories, denoted as ${\mathbf{z}}_{\text{w/cat}}(t)$ and ${\mathbf{z}}_{\text{w/o cat}}(t)$, using identical Gaussian noises and then compute the $1$-norm of the difference between these trajectories at each time point $t$ and sum them. We replicate the sampling $1000$ times independently and compute the mean, which represents the disparity between two samplers. As shown in Table 4, $\|{\mathbf{z}}_{\text{w/cat}}-{\mathbf{z}}_{\text{w/o cat}}\|_{1}$ gets smaller as the step size decreases.
[…]
NFE 10 12 20 30 40 50 100 200 1000
BFN 66.27 70.73 79.54 82.65 84.01 84.88 86.14 86.50 86.69
BFN w/o CS (ours) 70.67 73.63 80.86 83.26 84.27 84.90 86.11 86.76 86.77
Table 4: Ablation of the categorical sampling (CS) step in the BFN sampler on discrete data. We present the $L1$ norm of the difference between sampling trajectories of the original BFN sampler with and without CS, varying the step size.
Step size 0.1 0.05 0.02 0.01 0.005 0.002 0.001
$\|{\mathbf{z}}_{\text{w/cat}}-{\mathbf{z}}_{\text{w/o cat}}\|_{1}$ 0.18 0.16 0.13 0.10 0.087 0.066 0.049